# A Multi-Perspective Analysis of Memorization in Large Language Models

### v2

@@ Abstract
Memorization behavior of Large Language Models (LLMs) means LLMs can generate the same content used to train them. Though previous research has discussed memorization, the memorization of LLMs still lacks explanation, especially the cause of memorization and the dynamics of generating them. In this study, we comprehensively discussed memorization from various perspectives and extended the discussion to less and unmemorized content. Through various studies, we found that: (1) Through experiments, we revealed the relation of memorization between model size, continuation size, and context size. Further, we showed the dynamics of how unmemorized sentences transition to memorized sentences. (2) Through analysis of input, we found the frequency of tokens related to memorization and discovered the boundary effect when generating memorized or unmemorized sequences. (3) Through an analysis of embedding and decoding, we showed the LLM can generate paraphrased memorized sentences and sentences of different memorization scores clustered in the embedding space. Additionally, the entropy also shows a boundary effect when generating memorized or unmemorized content. (4)We trained a Transformer model to predict the memorization of different models, showing the possibility of predicting memorization and paraphrase memorization. Further analysis shows that the prediction of unmemorized sequences is easier than the prediction of memorized content, which can be explained by the significance of the boundary effect.
@@ 1 Introduction
  • In both the n-gram and decoding entropy dynamics, we found boundary phenomenon exists when the model generates memorized or unmemorized sentences. In the n-gram, the frequency when generating unmemorized tokens suddenly decreases while increasing for memorized tokens. In the entropy analysis, the entropy suddenly increases when generation unmemorized sentences while decreasing for memorized sentences.
@@ 3.1.1 Memorization Criteria
In the above equation, $X=\{x_{0}\cdots x_{n}\}$ is the sequence of predicted continuation tokens, and $Y=\{y_{0}\cdots y_{n}\}$ is the true continuation tokens. $L(Y)$ means the length of the true continuation tokens. If $M(X,Y)==1$, this sequence is completely memorized under the context sequence, which called as K-extractable. Similarly, we also define the sequence $Y$ as completely not memorized if $M(X,y)==0$. Since the memorization score has high granularity, in some experiments, we also classify sentences into several memorization levels based on their memorization scores.
@@ 3.2 Model Setting
Specifically, we investigate the model size of [70m, 160m, 410m, 1b, 2.8b, 6.9b, 12b] where $m$ stands for million and $b$ stands for billion. We choose the model trained on the deduplicated pre-train data version to avoid the effect of duplicated sentences, as previous research reports that the chance to be memorized grows exponentially with the number of duplicates. [^1] For more details, please refer to https://github.com/EleutherAI/pythia
@@ 4.1 Memorization Factors
This section discusses how memorization is affected by different factors. We collect the number of sentences memorized across different extend under different model sizes. Then, we divide them into 10 sets with a memorization difference of 0.1 The results are plotted in Figure 2.
@@ 4.1.1 The Factor of Model Size
  • We can see that sentences with low memorization scores (0-0.3) are significantly higher than sentences with high memorization scores. This shows that even though we can find memorization in the LLM, most of the contents are still not memorized.
  • Among sentences with high memorization scores, the memorized sentence increases faster than others. This indicates that LLM tends to memorize sentences fully rather than memorizing a piece of it. We can also see that the number of sentences with low memorization scores, like 0 and 0.1, decreases with the increase of model size, showing that the LLM tends to memorize the corpora with the increase of model.
  • Additionally, we can see that the sentences that are being memorized or memorized do not increase or decrease linearly with the expansion in model size. For example, we can see that for fully memorized sentences, the increase in sentence numbers is obvious from 70m to 2.8b compared to the increase from 2.8b to 12b. Similarly, the same decreasing trend can be seen from this figure for the unmemorized sentence. This shows that even with enough model size, the LLM may still be unable to memorize the whole corpora.
@@ 4.1.2 The Factor of Context Size and Complement Size
  • The decrease in memorized sentences with the complement size is not linear. Especially we can see that some sentences are firmly memorized, and even when we increase the complement size from 64 to 96, the number of memorized sentences does not change too much compared to the change when the complement size changes from 32 to 48.
  • We can also see that the drop in the number of memorized sentences with the complement size is more significant in the larger model. This suggests that the increase in model size significantly helps to memorize short sequences rather than long sequences.
  • The increase of memorized sentences with the context size is also not linear. The increase in context leads to a near-exponential increase in the number of memorized sentences. This suggests that with more information in the context, the more LLM can generate corresponding continuations.
@@ 4.2 Memorization Transition Across Model Size
This section discusses how sentences with different memorization scores transit across model sizes. Specifically, when trained with a larger model, we study the mutual transition between sentences with different memorization scores. We assigned sentences with memorization scores in the range of 0.2 with labels of very low, low, medium, high, and very high. We plot the transition matrix in Figure 3, which shows the transition from the 410m size model to the 2.8b size and 2.8b size to the 12b size model.
From the Figure 3, we can conclude the following:
@@ 4.3 Input Dyanamics
In this section, we discuss the question of whether there is any sign when the model starts to generate memorized or unmemorized content. Especially what makes sentences memorized at different states different and why some sentences are memorized by large-size models but not small-size models.
@@ 4.3.1 Token Level Frequency Analysis
Firstly, we begin with the input level, which is the pre-train corpora. We conduct this study by analyzing the n-gram statistics. First, we show the 1-gram statistics across steps in Figure 4, and we can see:
@@
  • The positive boundary effect of a memorized sentence suggests that the memorized sentence is caused by the first several tokens having a higher frequency, showing the frequency of the initial tokens is more important than the following tokens, e.g., remember the first several tokens make the following sentence to be memorized more easily. The negative boundary effect of the unmemorized sentences indicates the same results that the low frequency of the first several tokens makes the following sentence also easy to forget.
@@ 4.3.2 Sentence Level Frequency Analysis
  • First, we can see that the frequency of one-gram is significantly higher than the frequency of two-gram. In the context and continuation, the average frequency is roughly 3.5 times higher than that of the two-gram frequency. Additionally, such boundary effect is consistent in the two-gram setting. However, the positive boundary effect is less obvious due to the decrease of frequency when computed by 2-gram. However, as the unit is billion, the actual frequency gap is still a million level.
@@ 4.4.1 Embedding Dynamics
Figure 5: Embedding Dynamics of Across Different Model Size. Memorized Token Num $x$ means $x$ tokens generated align with the true continuation.
From the Figure 5, we can obtain the following points:
@@
  • Additionally, we can see that for highly memorized sentences, even when they are not totally memorized. They are still close in the embedding space indicating that the generated content may be both semantic and lexically similar to the memorized content.
  • Additionally, we can see that the large-size model has a higher Euclidean distance and lower Consine Similarity. The high Euclidean distance is caused by the expansion of hidden size (512 for 70m, 2048 for 1b, etc.). As for the decrease of the cosine similarity, this can also be attributed to the hidden size as the increase in hidden size expands the embedding space. This gives more expressivity to the embedding space, as the Euclidean distance between sentence embedding also expands, which leads to the decrease of Cosine Similarity. This may also help to explain the performance gap between different model sizes. Large-size models have more expressivity in the embedding space; thus, different sentences with different meanings are more clearly distributed with fewer neighbors in the space. The smaller embedding space in a small model makes the mix of embedding easier, leading to ambiguity in the embedding space, thus degrading the performance.
@@ 4.4.2 Generation Dynamics and Entropy
  • Additionally, when the model starts to generate unmemorized tokens, we can see a significant increase in entropy, showing such boundary effect also exists in the decoding steps while behaving oppositely compared to the boundary effect in frequency analysis. For the memorized sequences, we can also see a drop in entropy when decoding them. This corresponds to the whole memorized tokens and the first half of the half-memorized sequences; the entropy drops around the index of 32, which is the first generated token. However, we can see the drop of entropy for the memorized sentence is not samely significant as the half-memorized sentence, showing that even for the memorized content, the model is more confident when the sentence is fully memorized, leading to lower entropy.
@@
  • We can also see the entropy decrease with the increase in model size. This suggests that the larger size model is more confident about the generation of tokens compared to the small size model. Additionally, we can even see that the entropy is also lower in the context of memorized tokens, showing that the entropy of context also relates to how a sentence is memorized.
@@ 4.5 Prediction of Memorization
In this section, we train a Transformer model to predict the memorization of tokens from the LLM’s embedding and statistic features and analyze the results. We discuss whether Transformer is capable of this task and the difficulty of predicting the memorized token at different model sizes. As in the real world and indicated by our experiment results in previous sections, similar context sentences can also trigger memorized texts Stoehr et al. (2024), and the generated token can also be a paraphrase of the actual continuation Ippolito et al. (2023).
@@ 4.5.1 Results on Memorization of Prediction
  • First, regarding Token-level accuracy, we can see with a naive Transformer model, the token-level accuracy can reach $80\%$ and even higher, showing that the prediction of memorization at the token level is easy. The Full-level accuracy is low as this requires a correct prediction at every token.
@@
  • Thirdly, we found that with the increase in continuation size, the Token level accuracy gradually increases, possibly due to the increase in training data. For example, if the length to be predicted is $64$, the token used to train the model is four times higher than the prediction length of $16$. However, we see a drop in the Full accuracy as the increase in prediction length leads to more tokens being predicted. This makes reaching Full accuracy harder as the tokens required predicted grow.
@@ 4.5.2 Analysis of Full Accuracy
  • For any model size, most sentences that are fully predicted share low memorization scores. Especially, the 00 score (unmemorized) sentence occupies the largest portion among fully predicted sentences. Meanwhile, we can see that sentences with high memorization scores are hard to fully predict. We can see that the $1$ score (memorized) sentence has the lowest share among fully predicted sentences, showing that memorized sentences are hard to predict.
@@
  • The reason why a sentence with a low memorization score can be easily predicted can be due to the difference of significance at the boundary. As obtained from our previous experiments, the boundary effect of unmemorized sequences is more significant compared to memorized sequences in either token frequency or entropy of the generated token. Especially we can see that the boundary effect of memorized sequences decreases with model size while increasing for unmemorized sequences in previous results. This helps to explain why predicting memorization is easier in small-size models but harder in large-size models because the boundary effect becomes insignificant when we increase the model size. Similarly, the increase of boundary effect for unmemorized sentences with model size makes them easier to predict.
@@ 4.6 Conclusion
In this research, we comprehensively discussed the memorization behavior of LLMs across different perspectives. From a statistical level, we analyzed the memorization behavior of LLMs and expanded the analysis to sentences with lower memorization scores. Further transition analysis revealed the transition dynamics when trained in a large model, showing that most sentences stay in the original state while the memorized state is hard to transition to the unmemorized state. In the input level, through frequency analysis, we found the positive and negative boundary effect when generating memorized and unmemorized tokens and showed the LLM tends to memorize longer sequences rather than short pieces. In the output decoding level, in the embedding level, we found sentences with different memorization scores clusters in the embedding space, and the close distance of sentences with high memorization scores indicates the existence of paraphrase memorization. At the decoding dynamics, we conducted an analysis at the distribution level and calculated the entropy. We found that when the LLM starts to generate an unmemorized sentence, the entropy boundary effect also exists and significantly changes with the model size and the memorization score of the corresponding sentence. Finally, we trained a Transformer model to predict the memorization. We found that prediction at the token level is easy, while prediction at the sentence level is challenging. Through analysis of correctly fully predicted samples, we found unmemorized tokens are easy to predict while memorized tokens are hard to predict, and such a phenomenon can be explained by the significance of the boundary effect.
@@ 5 Limitations
This research has analyzed various factors regarding the memorization phenomena of LLMs. We acknowledge that there are still limitations to this research. Due to the lack of LLMs whose models and data are both being released, as the data is also different, it is hard to compare the memorization across different LLMs. How to comprehensively measure memorization of various LLMs, either close-sourced like GPT-4 or open-sourced like Pythia or LLaMa, can be a future research. Additionally, Pythia only provides LLMs up to 12B, which is still smaller than SotA open-sourced LLMs like LLaMA, whose largest model is 70B. The difference between 12B and 70B may lead to emergent abilities, which may also affect the memorization results obtained in 12B, which may not be able to scale even further due to those factors. Though we trained the Transformer to predict the memorization of LLMs, it is more analysis-oriented; thus, the performance is not the main focus of this experiment. Additionally, in this research, the discussion of memorization is under verbatim memorization, where generated tokens are identical to the same sentence in the corpora.
Appendix A Example Appendix
Figure 8: Prediction Examples
This is an appendix.

### v3

@@ Abstract
The memorization behavior of Large Language Models (LLMs) means that LLMs can generate the same content used to train them. Though previous research has discussed memorization, the memorization of LLMs still lacks explanation, especially the cause of memorization and the dynamics of generating them. In this study, we comprehensively discussed memorization from various perspectives and extended the discussion to less and unmemorized content. Through various studies, we found that: (1) Through experiments, we revealed the relation of memorization between model size, continuation size, and context size. Further, we showed the dynamics of how unmemorized sentences transition to memorized sentences. (2) Through input analysis, we found the frequency of tokens related to memorization and discovered the boundary effect when generating memorized or unmemorized sequences. (3) Through an analysis of embedding and decoding, we showed the LLM can generate paraphrased memorized sentences and sentences of different memorization scores clustered in the embedding space. The entropy also shows a boundary effect when generating memorized or unmemorized content. (4)We trained a Transformer model to predict the memorization of different models, showing the possibility of predicting memorization and paraphrase memorization. Further analysis shows that the prediction of unmemorized sequences is easier than the prediction of memorized content, which can be explained by the significance of the boundary effect.
@@ 1 Introduction
  • In both the n-gram and decoding entropy dynamics, we found that a boundary phenomenon exists when the model generates memorized or unmemorized sentences. In the n-gram, the frequency when generating unmemorized tokens decreases while increasing for memorized tokens. In the entropy analysis, the entropy suddenly increases when unmemorized sentences are generated, while it decreases for memorized sentences.
@@ 3.1.1 Memorization Criteria
In the above equation, $X=\{x_{0}\cdots x_{n}\}$ is the sequence of predicted continuation tokens, and $Y=\{y_{0}\cdots y_{n}\}$ is the true continuation tokens. $L(Y)$ means the length of the true continuation tokens. If $M(X,Y)==1$, this sequence is completely memorized under the context sequence, which is called K-extractable. Similarly, we also define the sequence $Y$ as completely not memorized if $M(X,y)==0$. Since the memorization score has high granularity, in some experiments, we also classify sentences into several memorization levels based on their memorization scores.
@@ 3.2 Model Setting
Specifically, we investigate the model size of [70m, 160m, 410m, 1b, 2.8b, 6.9b, 12b] where $m$ stands for million and $b$ stands for billion. We choose the model trained on the deduplicated pre-train data version to avoid the effect of duplicated sentences, as previous research reports that the chance to be memorized grows exponentially with the number of duplicates Kandpal et al. (2022). [^1] For more details, please refer to https://github.com/EleutherAI/pythia
@@ 4.1 Memorization Factors
This section discusses how memorization is affected by different factors. We collect the number of sentences memorized across different extend under different model sizes. Then, we divide them into ten sets with a memorization difference of 0.1 The results are plotted in Figure 2.
@@ 4.1.1 The Factor of Model Size
  • We can see that the number of sentences with low memorization scores (0-0.3) is significantly higher than sentences with high memorization scores. This shows that though memorization exists in the LLM, most of the pre-train data are still not memorized.
  • Among sentences with high memorization scores, the fully memorized sentence increases faster than others. This indicates that LLM tends to fully memorize sentences rather than memorizing a piece of it. We can also see that the number of sentences with low memorization scores (0, 0.1, etc) decreases with the increase in model size. This shows that unmemorized sentences will gradually memorized with the increase in model size.
  • Additionally, we can see that the sentences that are being memorized or memorized do not increase or decrease linearly with the increase in model size. For fully memorized sentences, the increase in sentence numbers is obvious from 70m to 2.8b compared to the increase from 2.8b to 12b. Similarly, the same decreasing trend can be seen from this figure for the unmemorized sentence. This shows an upper limit exists for memorization that the LLM cannot memorize the whole corpora even with enough model size.
@@ 4.1.2 The Factor of Context Size and Complement Size
  • The decrease in memorized sentences with the complement size is not linear. When we increase the complement size from 64 to 96, the number of memorized sentences does not change too much compared to the change when the complement size changes from 32 to 48. This also suggests some sentences are firmly memorized.
  • Additionally, the drop in the number of memorized sentences with the increased complement size is more significant in the larger model. This shows that though larger models memorize more sentences, the memorized sentence is not as firm as sentences memorized by small models.
  • The increase of memorized sentences with the context size is also not linear. The increase in context leads to a near-exponential increase in the number of memorized sentences. Additionally, the increase is more significant in the larger model, showing that more potential memorized contents can be elicited by giving more context compared to the small model.
@@ 4.2 Memorization Transition Across Model Size
This section discusses how sentences with different memorization scores transit across model sizes. Specifically, when trained with a larger model, we study the mutual transition between sentences with different memorization scores. We classify sentences with memorization scores with a range difference of 0.2 with labels of very low, low, medium, high, and very high. We plot the transition matrix in Figure 3, which shows the transition from the 410m size model to the 2.8b size and 2.8b size to the 12b size model.
From the Figure 3, we can conclude:
@@ 4.3 Input Dyanamics
In this section, we discuss the question of whether there is any sign when the model starts to generate memorized or unmemorized content. Especially what makes sentences memorized at different extents and why some sentences are memorized by large-size models but not small-size models.
@@ 4.3.1 Token Level Frequency Analysis
Firstly, we begin with the input level by analyzing the n-gram statistics of the pre-train corpora. We show the 1-gram statistics across steps in Figure 4, and we can see:
@@
  • The positive boundary effect of a memorized sentence suggests that the memorized sentence is caused by the first several tokens having a higher frequency, showing the frequency of the initial tokens is more important than the following tokens, e.g., remember the first several tokens make the following sentence to be memorized more easily. The negative boundary effect of the unmemorized sentences indicates the same results: the low frequency of the first several tokens makes the following sentence easy to forget.
@@ 4.3.2 Sentence Level Frequency Analysis
  • First, we can see that the frequency of one-gram is significantly higher than the frequency of two-gram. In the context and continuation, the average frequency is roughly 3.5 times higher than the two-gram frequency. Additionally, such boundary effect is consistent in the two-gram setting. However, the positive boundary effect is less obvious due to the decrease of frequency when computed by 2-gram. However, as the unit is billion, the actual frequency gap is still a million level.
@@ 4.4.1 Embedding Dynamics
Figure 5: Embedding Dynamics of Across Different Model Size. Memorized Token Num $x$ means $x$ tokens generated to align with the true continuation.
From Figure 5, we can obtain the following points [^2] Detailed numbers of Cosine Similarity and Euclidean distance between sentences with different memorization scores are at different decoding steps are at Appendix B.2:
@@
  • Additionally, we can see that for highly memorized sentences, even when they are not totally memorized. They are still close in the embedding space, indicating that the generated content may be both semantically and lexically similar to the memorized content.
  • Additionally, we can see that the large-size model has a higher Euclidean distance and lower Consine Similarity. The high Euclidean distance is caused by the expansion of hidden size (512 for 70m, 2048 for 1b, etc.). The decrease of the cosine similarity can also be attributed to the hidden size, as the increase in hidden size expands the embedding space. This gives more expressivity to the embedding space, as the Euclidean distance between sentence embedding also expands, which leads to the decrease of Cosine Similarity. This may also help to explain the performance gap between different model sizes. Large-size models have more expressivity in the embedding space; thus, different sentences with different meanings are more clearly distributed with fewer neighbors in the space. The smaller embedding space in a small model makes the mix of embedding easier, leading to ambiguity in the embedding space, thus degrading the performance.
@@ 4.4.2 Generation Dynamics and Entropy
  • Additionally, when the model starts to generate unmemorized tokens, we can see a significant increase in entropy, showing such boundary effect also exists in the decoding steps while behaving oppositely compared to the boundary effect in frequency analysis. We can also see a drop in entropy when decoding the memorized sequences. This corresponds to the fully memorized tokens and the first half of the half-memorized sequences. The entropy drops around the index of 32, which is the first generated token. However, we can see the drop of entropy for the memorized sentence is not samely significant as the half-memorized sentence, showing that even for the memorized content, the model is more confident when the sentence is fully memorized, leading to lower entropy.
@@
  • We can also see the entropy decrease with the increase in model size. This suggests that larger models are more confident about generating tokens than small ones. Additionally, we can see that the entropy is lower in the context of memorized tokens, showing that the entropy of context also relates to how a sentence is memorized.
@@ 4.5 Prediction of Memorization
In this section, we train a Transformer model to predict the memorization of tokens from the LLM’s embedding and statistic features and analyze the results. We discuss whether Transformer can do this task and the difficulty of predicting the memorized token at different model sizes. As in the real world and indicated by our experiment results in previous sections, similar context sentences can also trigger memorized texts Stoehr et al. (2024), and the generated token can also be a paraphrase of the actual continuation Ippolito et al. (2023).
@@ 4.5.1 Results on Memorization of Prediction
  • First, regarding Token-level accuracy, we can see with a naive Transformer model, the token-level accuracy can reach $80\%$ and even higher, showing that the prediction of memorization at the token level is easy. The full accuracy is low as this requires a correct prediction for every token.
@@
  • Thirdly, we found that with the increase in continuation size, the Token level accuracy gradually increases, possibly due to the increase in training data. For example, if the length to be predicted is $64$, the token used to train the model is four times higher than the prediction length of $16$. However, we see a drop in the Full accuracy as the increase in prediction length leads to more tokens being predicted. This makes reaching full accuracy harder as the tokens required to be predicted grow.
@@ 4.5.2 Analysis of Full Accuracy
  • For any model size, most sentences that are fully predicted share low memorization scores. The 00 score (unmemorized) sentence occupies the largest portion of fully predicted sentences. Meanwhile, we can see that sentences with high memorization scores are hard to predict fully and correctly. We can see that the $1$ score (memorized) sentence has the lowest share among fully predicted sentences, showing that memorized sentences are hard to predict.
@@
  • The reason why a sentence with a low memorization score can be easily predicted can be due to the difference of significance at the boundary. As obtained from our previous experiments, the boundary effect of unmemorized sequences is more significant than memorized sequences in either token frequency or entropy of the generated token. Especially we can see that the boundary effect of memorized sequences decreases with model size while increasing for unmemorized sequences in previous results. This helps to explain why predicting memorization is easier in small-size models but harder in large-size models because the boundary effect becomes insignificant when we increase the model size. Similarly, the increase of boundary effect for unmemorized sentences with model size makes them easier to predict.
@@ 4.6 Conclusion
In this research, we comprehensively discussed the memorization behavior of LLMs across different perspectives. From a statistical level, we analyzed the memorization behavior of LLMs and expanded the analysis to sentences with lower memorization scores. Further transition analysis revealed the transition dynamics when trained in a large model, showing that most sentences stay in the original state while the memorized state is hard to transition to the unmemorized state. At the input level, through frequency analysis, we found the positive and negative boundary effects when generating memorized and unmemorized tokens. We showed the LLM tends to memorize longer sequences rather than short pieces. In the output decoding level, in the embedding level, we found clusters of sentences with different memorization scores in the embedding space, and the close distance of sentences with high memorization scores indicates the existence of paraphrase memorization. We analyzed the distribution level and calculated the entropy at the decoding dynamics. We found that when the LLM starts to generate an unmemorized sentence, the entropy boundary effect also exists and significantly changes with the model size and the memorization score of the corresponding sentence. Finally, we trained a Transformer model to predict the memorization. We found that prediction at the token level is easy, while prediction at the sentence level is challenging. Through analysis of correctly fully predicted samples, we found unmemorized tokens are easy to predict while memorized tokens are hard to predict, and such a phenomenon can be explained by the significance of the boundary effect.
@@ 5 Limitations
This research has analyzed various factors regarding the memorization phenomena of LLMs. We acknowledge that there are still limitations to this research. Due to the lack of LLMs whose models and data are both being released, as the data is also different, it is hard to compare the memorization across different LLMs. Future research can be on how to comprehensively measure memorization of various LLMs, either close-sourced like GPT-4 or open-sourced like Pythia or LLaMa. Additionally, Pythia only provides LLMs up to 12B, which is still smaller than SotA open-sourced LLMs like LLaMA, whose largest model is 70B. The difference between 12B and 70B may lead to emergent abilities, which may also affect the memorization results obtained in 12B, which may not be able to scale even further due to those factors. Though we trained the Transformer to predict the memorization of LLMs, it is more analysis-oriented; thus, the performance is not the main focus of this experiment. Additionally, in this research, the discussion of memorization is under verbatim memorization, where generated tokens are identical to the same sentence in the corpora.
Appendix A Appendix
Figure 8: Prediction Examples. Pred Prob means the output predicted probability of the corresponding label in the pred row for each example. Gold means the true label, Pred means the predicted label, and Pred Prob means the probability of the corresponding prediction. $M$ means the label for a token at this index is a memorized token. $U$ means the label for a token at this index is an unmemorized token.
Appendix B Experiment Setting
The experiment uses 64 A100 40Gbs GPUs when using LLMs to generate tokens given the previous context, which utilizes PyTorch’s parallel running packages. We run the model with half-precision, which increases both the speed and saves memory. This follows the previous Pythia implementation when generating tokens given context.
The running time depends on the model size and the generated token length. In the situation when a 70m model with 32 context tokens and 16 tokens is required to be generated, it can be run with one A100 GPU within several hours. However, if such an experiment is in a single A100 GPU, it would estimated to take two weeks to finish the generation. Therefore, with 64 GPUs, the running of the 12b model in such a situation can be shortened to around one day. However, the generation time also largely increases with the length of generated tokens, which grows in a linear way with the number of tokens to be generated. Additionally, since we use greedy decoding and do not consider other possible decoding options but the token with the highest probability, it is also possible that the running time may be different if using a more complicated decoding strategy. However, using a more complex decoding strategy does not affect the results. If a sequence of tokens is memorized, the memorized token will always be the most probable token when generating them.
B.1 Case Study
We also provide a case study of the model’s prediction on the test set regarding the prediction of memorization in Figure 8. From this figure, we can see that:
  1. 1.
    Confirming previous experiments, the memorized token is mostly continuous, showing the memorization happens in chunks of sequences rather than individual sequences.
  2. 2.
    In the first example, the model outputs a fully correct prediction that aligns with the actual label, showing the possibility of predicting memorization by utilizing embedding information.
  3. 3.
    In the second example, we see that the model’s prediction does not align with the actual continuation. The model predicts an unmemorization label for the memorized label.
  4. 4.
    In the third example, we can see this unmemorized sequence. The model fully predicted those labels. We also noticed that the probability for the corresponding token is very high, showing that the model is very confident about the prediction.
Figure 9: 410m Model, Step 2
Figure 10: 410m Model, Step 9
Figure 11: 410m Model, Step 16
Figure 12: 2.8b Model, Step 2
Figure 13: 2.8b Model, Step 9
Figure 14: 2.8b Model, Step 16
Figure 15: 12b Model, Step 2
Figure 16: 12b Model, Step 9
Figure 17: 12b Model, Step 16
B.2 Supplementary Data for Embedding Dynamics
We have presented the PCA visualized embedding dynamics in Figure 5. In this section, we provide the actual numbers of both Cosine Similarity and Euclidean distance to further illustrate this point.
Similarly to Figure 5, we also provide detailed cosine similarity and distance results from Figure 9 to Figure 17. From those figures, we can see that the cosine similarity fluctuates but remains relatively stable for sentences with different memorization scores. Additionally, sentences that are not exactly the same but close regarding the memorization score are very close in the embedding space. For example, the fully memorized sentence is also close to sentences with high memorization scores with embedding similarity over 0.9. This shows the possibility that the model is generating paraphrased memorized sequences. However, with the increase in the model size, the cosine similarity decreases. For the Euclidean distance, we can see that the embedding distances have a decreasing trend with the increasing decoding steps, while the mutual Euclidean increases with the model size.
Figure 18: Index Distribution of Memorized Sequences in 12b model.
B.3 Does LLM prefer to memorize specific parts within the training data?
In this section, we discuss the question of whether LLM prefers to memorize a specific part within the training data. We split the corpora into 50 parts based on their index and examine how many memorized sentences are in those parts.
From the result shown in Figure 18, though the number of memorized sentences is not completely evenly distributed, we can see that there is no significant part that the number of memorized sentences is clearly more than others. This shows the training order does not affect memorization.